# Paravortices: loop braid representations with both generators involutive

### v1

@@ 1. Introduction
The study of local representations of the loop braid groups $L_{n}$ was initiated in [KMRW], partially motivated by the desire to explore topological phases of matter in 3 spatial dimensions. The motions of free loops (e.g., vortices) in 3 dimensional media are natural generalisations of the motions of points (e.g., anyons) in 2 dimensions, so $L_{n}$ plays the role of the braid group $B_{n}$ in the higher dimensional sequel to the topological quantum computation via anyons story [FLKW]. Indeed, $L_{n}$ resembles $B_{n}$ strongly: the motions of passing a loop through its neighbour (‘leapfrogging’) form a subgroup isomorphic to $B_{n}$, while interchanging adjacent loops have order 2 and yield a copy of the symmetric group $\Sigma_{n}$.
By a local representation we mean a strict monoidal functor $F$ from the loop braid category $\mathbf{L}$ to the category $\mathsf{Mat}$ of matrices, as in [MRT25] for the braid category $\mathsf{B}$ setting [MacLane]. The morphisms from object $n$ to itself in $\mathbf{L}$ form the loop braid group $L_{n}$, just as the morphisms in $\mathsf{B}$ are braid groups $B_{n}$. In practice, one specifies such a functor $F$ by means of two Yang-Baxter operators $R,S$ satisfying certain compatibility relations and $S^{2}=I$. There are numerous reasons to focus on these local representations, not least of which is the connection to the quantum circuit model for quantum computation, in which circuits are cascades of matrices acting on a few adjacent factors of $\bigotimes_{i}{\mathbb{C}}^{n}$.
In this article we study monoidal functors from the mixed doubles category obtained by setting also $R^{2}=1$ and denoted $\mathsf{MD}$}. The morphism groups in this category are thus quotients ${\mathsf{MD}}_{n}$ of $L_{n}$ by the relation that the leapfrogging generators also have order $2$. Thus ${\mathsf{MD}}_{n}$ has two copies of the symmetric group $\Sigma_{n}$ on equal footing, satisfying some mixed relations: hence the ‘mixed doubles’ moniker.
@@
A physical motivation is that, philosophically, these are a natural 3-dimensional analogue of paraparticles as in [WangHazzard2024Supplement], where the exchange statistics of $n$ identical paraparticles yield representations of the symmetric group $\Sigma_{n}$. From this point of view, representations of ${\mathsf{MD}}_{n}$ correspond to paravortices: loop-like excitations for which the loop leapfrog operation on adjacent loops squares to the identity. Thus ${\mathsf{MD}}_{n}$ fits into the topological quantum computation story as the group governing the loop motion statistics of paravortices. Representations of ${\mathsf{MD}}_{n}$ can then be studied as quantum gates implemented on systems of paravortices, which could provide fault-tolerant models for quantum computation. We will see that universality is unlikely, but they are a compelling potential quantum resource, nonetheless.
@@
Mathematically, ${\mathsf{MD}}_{n}$ is a non-trivial quotient of $L_{n}$, analogous to $\Sigma_{n}$ as the first non-trivial quotient of $B_{n}$. Thus ${\mathsf{MD}}_{n}$ is a Coxeter-style combinatorial version of $L_{n}$. Unlike $\Sigma_{n}$, ${\mathsf{MD}}_{n}$ is an infinite group, so it retains some of the intrigue of $L_{n}$. Moreover, the mixing of the two copies of $\Sigma_{n}$ has the effect of making the basis choices matter - a type of higher (or extended) representation theory.
@@
Next, although perhaps the main problem in this area is the classification of braid representations, the part of it corresponding to the classification of symmetric representations (strict monoidal functors $F:\Sigma\rightarrow\mathsf{Mat}$) is also a current focus. And of course $\mathsf{MD}$} representations can be seen as a generalisation of this. In the symmetric rep context, there is a classification of unitary, involutive solutions up to a representation-theoretic notion of equivalence due to [LechnerPennigWood]. Thus the representation theory of $\mathsf{MD}$} potentially interpolates between the full braid representation problem - classifying strict monoidal functors $F:\mathsf{B}\rightarrow\mathsf{Mat}\;$ - and partial solutions such as the [LechnerPennigWood] classification of symmetric representations/involutive braid reps $F:\Sigma\rightarrow\mathsf{Mat}$.
An intriguing but relatively tame presentation of $L_{n}$ as a semidirect product is available [BardakovBellingeriDamiani15] which can be leveraged for representation theoretic computations.
@@
We also note that we can make contact with the growing community of researchers in set-theoretical solutions to the Yang-Baxter equation, where a main focus is on involutive solutions (see, e.g., [gatevaivanova2009multipermutation]). A natural subproblem of the classification of local ${\mathsf{MD}}_{n}$ representations would be to ask for pairs of set-theoretical solutions $(R,S)$ that yield such a representation. Contact between this area and the classification programme for braid representations is already discussed in [MRT25].
@@
Our main result, Theorem 4.5, is a classification, up to natural equivalences, of all strict monoidal functors $F:{\mathsf{MD}}\rightarrow\mathsf{Mat}$ with $F(1)=2$, i.e., all local representations with $R,S$ being $4\times 4$. In §5 we develop and deploy some machinery to enable us to analyse several of these functors as representations (i.e. as collections of ordinary representations taken together), providing, in some cases, all-$n$ descriptions of their irreducible content and their decompositions into indecomposable subrepresentations. In other notable cases we simply analyse representations in low rank by more direct methods.
@@ 2. Preliminaries
For $n\in\mathbb{N}$ the loop braid group $\mathbf{L}_{n}$ is a motion group of a configuration of $n$ unlinked unknotted oriented loops in 3-dimensions (see e.g. [Damiani] for a review and original references). For definiteness, for the reference configuration, we take circles of unit radius lying in the xy-plane, with centres regularly spaced along the x-axis. For example in $\mathbf{L}_{2}$ we can consider a certain pair of motions. Both involve the exchange of the two loops. Let us consider both motions in a frame in which the first loop appears static - so that the second loop appears to pass from a place to the right to a place to the left of it. In the first motion there is a representative concrete motion in which the moving loop completes its journey well away from the static loop. In the second motion it passes once through the disc of the static loop. It will be clear that these representatives give distinct motions; indeed that the first is involutive and that the second is not. In fact these two motions $\sigma$ and $\rho$ (say) generate $\mathbf{L}_{2}$, and a corresponding chain of pair exchange elements $\sigma_{i}$ and $\rho_{i}$ generates $\mathbf{L}_{n}$. Thus $\sigma_{i}^{2}=1$ and so on.
The group $\mathbf{L}_{n}$ is isomorphic to the abstract group with presentation
@@
The category $\mathbf{L}$ is the strict monoidal category with object class $\mathbb{N}_{0}$, with $\mathbf{L}(n,n)=\mathbf{L}_{n}$; $\mathbf{L}(n,m)=\emptyset$ if $n\neq m$; and monoidal product given by juxtaposition of the reference configurations along the x-axis (cf. the braid category in [MacLane], and see §2).
For $n\in\mathbb{N}$ the mixed doubles group ${\mathsf{MD}}_{n}$ is defined as the group
@@
Here we give several useful details of the representation theory of ${\mathsf{MD}}_{n}$, and of the collection of all these groups naturally arranged in a strict monoidal category - the mixed doubles category $\mathsf{MD}$}.
These groups rank-by-rank have various other names in the literature depending on the background of the authors. See e.g. [BardakovBellingeriDamiani15], where ${\mathsf{MD}}_{n}$ is called flat welded braids, and references therein.
The group ${\mathsf{MD}}_{n}$ is shown in [BardakovBellingeriDamiani15] to be isomorphic to a certain semidirect product ${\mathbb{Z}}^{\binom{n}{2}}\rtimes\Sigma_{n}$ (see below for details). In particular ${\mathsf{MD}}_{n}$ is virtually abelian. Classical Clifford theory [isaacs1994] can therefore be used in principle to describe the irreducible representations of ${\mathsf{MD}}_{n}$. Our interest is mainly in local representations, i.e. those associated with a pair of Yang-Baxter operators $(R,S)$ yielding a strict monoidal functor from the category $\mathsf{MD}$} build from the groups ${\mathsf{MD}}_{n}$ to the category $\mathsf{Mat}$ of matrices. In particular, such representations can be non-semisimple, pushing us into the world of indecomposable, rather than irreducible, constituent representations. Even for ${\mathbb{Z}}^{\binom{n}{2}}$ the classification problem for indecomposable representations is wild, so local representations are not easily decomposed.
@@
(2.1) Here ‘local representation’ means in essence that we consider representations for a natural-number-indexed sequence of presented groups, and that in each presentation every generator is of the form $\alpha_{i}$ where $\alpha$ is taken from the same finite set, $i$ ranges over natural numbers, and every relation involving two non-adjacent $i$’s is a commutation.
Note the following:
@@
It is clear from these relations that for $n>1$ the automorphism group of ${\mathsf{MD}}_{n}$ has a subgroup isomorphic to ${\mathbb{Z}}/2{\mathbb{Z}}\times{\mathbb{Z}}/2{\mathbb{Z}}$ generated by $s_{i}\leftrightarrow r_{i}$ and the simultaneous $(s_{i},r_{i})\leftrightarrow(s_{n-i},r_{n-i})$.
@@ 3. Generalised Wreaths and the BaBeDa isomorphism
We will not need the following until §A, but it is convenient to establish now.
(3.2) Since prepending an automorphism takes a $B$-rep to a $B$-rep, we have an action of CC on the set of $B$-representations. In particular it takes irreducibles to irreducibles. But in general let us write $\rho^{c}$ for the image of $B$-rep $\rho$ under the action of $c\in C$.
@@ 3.1. Warm-up: wreaths, their conjugacy classes, and their representations
(3.4) Recall further that there is a natural ‘bird-track’ calculus (cf. e.g. [Cvitanovic, MartinWoodcockLevy00]) for $G\wr\Sigma_{n}$ which decorates the strings of the usual diagram calculus for $\Sigma_{n}$ with elements of $G$. Writing $\tau\in G$ for a generating element (and $\sigma_{1}\in\Sigma_{n}$ as usual), then the calculus looks like this:
(all such figures here are taken directly from [MartinWoodcockLevy00]). The calculus allows $\sigma_{1}\tau_{1}\sigma_{1}$ to be represented as a bead on the second track:
@@ 3.2. Generalised wreaths
(3.6) For $n\in\mathbb{N}$ consider the group ${\mathbb{Z}}^{{n}\choose{2}}$, with generators $x_{ij}$ for $1\leq i<j\leq n$. That is, ${\mathbb{Z}}^{{n}\choose{2}}=\;\langle x_{ij},i<j\;|x_{ij}x_{kl}=x_{kl}x_{ij}\rangle\;=\;\{\prod_{i<j}x_{ij}^{a_{ij}}\;|\;a_{ij}\in{\mathbb{Z}}\}$. We may write $g\in{\mathbb{Z}}^{{n}\choose{2}}$ as $(x_{12}^{a_{12}},x_{13}^{a_{13}},..)$. In particular that is $x_{12}=(x_{12}^{1},1,1,..,1)$ and $1=(1,1,..,1)$. …
@@
(3.8) In ${\mathsf{MD}^{\prime}}_{n}$ define $\varsigma_{i}$ as $(1,\sigma_{i})=((1,1,\ldots,1),\sigma_{i})$ (where $\sigma_{i}$ is the elementary transposition in $\Sigma_{n}$) and $x_{ij}=(x_{ij},1)$. Observe that ${\mathsf{MD}^{\prime}}_{n}$ admits a presentation with generators $x_{12}$ and $\varsigma_{i}$, $i=1,2,\ldots,n-1$ together with the usual symmetric group relations, the relation
@@ 3.3. The BaBeDa isomorphism
(3.9) Consider the map $\varphi$ on the generators of ${\mathsf{MD}^{\prime}}_{n}$ (and hence on the free group with these formal generators) to ${\mathsf{MD}}_{n}$ given by $\varsigma_{i}\mapsto s_{i}$ and $x_{12}\mapsto r_{1}s_{1}$. It is shown in [BardakovBellingeriDamiani15, Prop.5.5] that $\varphi$ extends to an isomorphism ${\mathsf{MD}^{\prime}}_{n}\cong{\mathsf{MD}}_{n}$, which we call the BaBeDa isomorphism as an portmanteau of the authors’ surnames. We also give the proof here to give a self-contained explanation of how the copy of ${\mathbb{Z}}^{{n}\choose{2}}$ arises, for which we need some preliminary results given here.
(3.10) The virtual braid group ${\mathsf{VB}}_{n}$ is obtained by taking the same generators as $L_{n}$, and taking as relations the leftmost and rightmost relation in (1), as well as all relations in (2) and (3). There is a map $\pi\colon VB_{n}\to\Sigma_{n}$ given by letting $r_{i}$ and $s_{i}$ map to the generator $(i,i+1)$. This is clearly surjective, thus defining ${\mathsf{VP}}_{n}$ to be the kernel of this map gives that ${\mathsf{VB}}_{n}$ breaks down as a semi direct product ${\mathsf{VP}}_{n}\rtimes\Sigma_{n}$. It is proved in [Bardakov, Th.1] that $VP_{n}$ admits a presentation with generators $\lambda_{kl}$, $1\leq k\neq l\leq n$, and relations
@@
$\lambda_{ki}\lambda_{kj}\lambda_{ij}=\lambda_{ij}\lambda_{kj}\lambda_{ki}.$ (14)
The group $VP_{n}$ is realised as a subgroup of $VB_{n}$ by letting
@@
(3.11) It is also shown in [Bardakov, Le.1] The subgroup $VP_{n}\subset{\mathsf{VB}}_{n}$ satisfies the following conjugation relation
@@
First note that it is clear that all relations involving only the symmetric group generators in each group are clearly preserved by both $\varphi$ and $\hat{\varphi}$.
@@
whenever $k\neq i$. This can be seeing by observing that, using the conjugation rule (3.3) and that $\varphi(x_{ij})=\lambda_{ij}$,

### v2

@@ 1. Introduction
The study of local representations of the loop braid groups $L_{n}$ was initiated in [16], partially motivated by the desire to explore topological phases of matter in 3 spatial dimensions, see also for example [26, 5]. The motions of free loops (e.g., vortices) in 3 dimensional media are natural generalisations of the motions of points (e.g., anyons) in 2 dimensions, so $L_{n}$ plays the role of the braid group $B_{n}$ in the higher dimensional sequel to the topological quantum computation via anyons story, see [12, 24, 28] and references therein. Indeed, $L_{n}$ resembles $B_{n}$ strongly: the motions of passing a loop through its neighbour (‘leapfrogging’) form a subgroup isomorphic to $B_{n}$, while interchanging adjacent loops have order 2 and yield a copy of the symmetric group $\Sigma_{n}$.
By a local representation we mean a strict monoidal functor $F:\mathbf{L}\rightarrow\mathsf{Mat}$ from the loop braid category $\mathbf{L}$ to the category $\mathsf{Mat}$ of matrices (see e.g. [22] for a review of the setup in the setting of the braid category $\mathsf{B}$ [18]). The morphisms from object $n$ to itself in $\mathbf{L}$ form the loop braid group $L_{n}$, just as the morphisms in $\mathsf{B}$ are braid groups $B_{n}$. In practice, one specifies such a functor $F$ by means of two Yang-Baxter operators $R,S$ satisfying certain compatibility relations and $S^{2}=I$. There are numerous reasons to focus on these local representations, not least of which is the connection to the quantum circuit model for quantum computation, in which circuits are cascades of matrices acting on a few adjacent factors of $\bigotimes_{i}{\mathbb{C}}^{n}$ [25]. For existing examples in the literature of local loop braid representations in physical settings see also for example [5, 8] and references therein.
In this article we study monoidal functors from the mixed doubles category, which is obtained by setting also $R^{2}=1$ and is denoted $\mathsf{MD}$}. The morphism groups in this category are thus quotients ${\mathsf{MD}}_{n}$ of $L_{n}$ by the relation that the leapfrogging generators also have order $2$. Thus ${\mathsf{MD}}_{n}$ has two copies of the symmetric group $\Sigma_{n}$ on equal footing, satisfying some mixed relations: hence the ‘mixed doubles’ moniker.
@@
A physical motivation is that, philosophically, these are a natural 3-dimensional analogue of paraparticles [30] (a generalisation of parafermions, see eg. [1]) where the exchange statistics of $n$ identical paraparticles yield representations of the symmetric group $\Sigma_{n}$. From this point of view, representations of ${\mathsf{MD}}_{n}$ correspond to paravortices: loop-like excitations for which the loop leapfrog operation on adjacent loops squares to the identity. Thus ${\mathsf{MD}}_{n}$ fits into the topological quantum computation story as the group governing the loop motion statistics of paravortices. Representations of ${\mathsf{MD}}_{n}$ can then be studied as quantum gates implemented on systems of paravortices, which could provide fault-tolerant models for quantum computation. We will see that universality is unlikely, but they are a compelling potential quantum resource, nonetheless (cf. [27]).
@@
Mathematically, ${\mathsf{MD}}_{n}$ is a non-trivial quotient of $L_{n}$, analogous to $\Sigma_{n}$ as the first non-trivial quotient of $B_{n}$. Thus ${\mathsf{MD}}_{n}$ is a Coxeter-style combinatorial version of $L_{n}$. Unlike $\Sigma_{n}$, ${\mathsf{MD}}_{n}$ is an infinite group, so it retains some of the intrigue of $L_{n}$. Moreover, the mixing of the two components - the two copies of $\Sigma_{n}$ - has the effect of making basis choices matter, which can be cast as a type of higher (or extended) representation theory problem. An intriguing but relatively tame presentation of $L_{n}$ as a semidirect product is available [4] which can be leveraged for representation theoretic computations, as we shall discuss.
@@
Next, although perhaps the main problem in this area is the classification of braid representations, the part of it corresponding to the classification of symmetric representations (strict monoidal functors $F:\Sigma\rightarrow\mathsf{Mat}$) is also a current focus. And of course $\mathsf{MD}$} representations can be seen as a generalisation of this. In the symmetric rep context, there is a classification of unitary, involutive solutions up to a representation-theoretic notion of equivalence, due to [17]. Thus the representation theory of $\mathsf{MD}$} potentially interpolates between the full braid representation problem - classifying strict monoidal functors $F:\mathsf{B}\rightarrow\mathsf{Mat}\;$ - and partial solutions such as the [17] classification of symmetric representations/involutive braid representations $F:\Sigma\rightarrow\mathsf{Mat}$.
@@
We also note that we can make contact with the growing community of researchers in set-theoretical solutions to the Yang-Baxter equation, where a main focus is on involutive solutions (see, e.g., [13]). A natural subproblem of the classification of local ${\mathsf{MD}}_{n}$ representations would be to ask for pairs of set-theoretical solutions $(R,S)$ that yield such a representation. Contact between this area and the classification programme for braid representations is already discussed in [22].
@@
Our main result, Theorem 4.5, is a classification, up to natural equivalences, of all strict monoidal functors $F:{\mathsf{MD}}\rightarrow\mathsf{Mat}$ in rank 2, i.e. with $F(1)=2$. That is, we classify all local representations with $R,S$ being $4\times 4$. In §5 we develop and deploy some machinery to enable us to analyse several of these functors as representations (i.e. as collections of ordinary representations taken together), providing, in some cases, all-$n$ descriptions of their irreducible content and their decompositions into indecomposable subrepresentations. In other notable cases we simply analyse representations in low rank by more direct methods.
Acknowledgements. PM thanks EPSRC for funding under project EP/W007509/1; and thanks the Royal Society for funding ER’s Wolfson Fellowship at Leeds contributing to this work. PM also thanks Jarmo Hietarinta, Sarah Almateari and Paula Martin for useful discussions. The work of ER was also partially supported by US NSF grant DMS-2205962. Part of this work was carried out while ER was visiting the University of Leeds on a Royal Society Wolfson Fellowship, and he thanks them for their hospitality and support. And all 3 authors thank Zhenghan Wang for continued inspiration and guidance.
@@ 2. Preliminaries
For $n\in\mathbb{N}$ the loop braid group $\mathbf{L}_{n}$ is a motion group of a configuration of $n$ unlinked unknotted oriented loops in 3-dimensions (see e.g. [9] for a review and original references). For definiteness, for the reference configuration, $C_{n}$, we take circles of unit radius lying in the xy-plane, with centres regularly spaced along the x-axis. Hence in particular for each loop $l$ in the reference configuration there is a disk of which $l$ is the boundary. For example in $\mathbf{L}_{2}$ we can consider a certain pair of motions. Both involve the exchange of the two loops. Let us consider both motions in a frame in which the first loop appears static - so that the second loop appears to pass from a place to the right to a place to the left of it. (N.B., in this frame the constant representative of the identity motion has neither loop apparently static to the observer.) In the first motion, $\sigma$, say, there is a representative concrete motion in which the moving loop completes its journey well away from the static loop. In the second motion, $\rho$, there is a representative in which the moving loop passes once through the disc of the static loop. It will be clear that these representatives give distinct motions; indeed that the first is involutive and that the second is not.
In fact the two motions $\sigma$ and $\rho$ generate $\mathbf{L}_{2}$, and, fixing $n$, a corresponding chain of pair exchange elements $\sigma_{i}$ and $\rho_{i}$ generates $\mathbf{L}_{n}$. Thus $\sigma_{i}^{2}=1$ and $\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i}$ for $j\neq i\pm 1$, and so on.
(2.1) The category $\mathbf{L}$ is the strict monoidal category with object class $\mathbb{N}_{0}$, with $\mathbf{L}(n,n)=\mathbf{L}_{n}$; $\mathbf{L}(n,m)=\emptyset$ if $n\neq m$; and monoidal product given by juxtaposition of the reference configurations along the x-axis (see for example [19] and references therein, and cf. the braid category in [18]).
Observe that category $\mathbf{L}$ is generated as a monoidal category by the object 1 and the morphisms $\sigma$ and $\rho\in\mathbf{L}(2,2)$. In particular , writing $1_{1}$ for the identity in $\mathbf{L}(1,1)$ we have (fixing $n=3$ for a moment) $\sigma_{2}=1_{1}\otimes\sigma\in\mathbf{L}(3,3)$, and so on.
(2.2) Consider the abstract group $L_{n}$ with presentation
@@
(2.3) For each $n$ the group $\mathbf{L}_{n}$ is isomorphic to the abstract group $L_{n}$ under the map on generators given by $\sigma_{i}\mapsto s_{i}$ and $\rho_{i}\mapsto r_{i}$.
(2.4) For $n\in\mathbb{N}$ the mixed doubles group ${\mathsf{MD}}_{n}$ is defined by the presentation
@@
The groups ${\mathsf{MD}}_{n}$ rank-by-rank have various other names in the literature depending on the background of the authors. See e.g. [4], where ${\mathsf{MD}}_{n}$ is called the group of flat welded braids, and references therein.
@@
(2.7) The collection of all the groups ${\mathsf{MD}}_{n}$ naturally arranged in a strict monoidal category is the mixed doubles category $\mathsf{MD}$}.
A monoidal functor $F:{\mathsf{MD}}\rightarrow\mathsf{Mat}$ is here called a mixed doubles representation (following the nomenclature of spin-chain braid representation, shortened to braid representation, in [20]).
The group ${\mathsf{MD}}_{n}$ is shown in [4] to be isomorphic to a certain semidirect product ${\mathbb{Z}}^{\binom{n}{2}}\rtimes\Sigma_{n}$ (see §3 for details). In particular ${\mathsf{MD}}_{n}$ is virtually abelian. Classical Clifford theory [15] can therefore be used in principle to describe the irreducible representations of ${\mathsf{MD}}_{n}$. In practice this approach is not straightforward.
Our interest here is in mixed doubles representations, as above. These are examples of local representations, i.e. representations associated with a pair of Yang-Baxter operators $(R,S)$ yielding a strict monoidal functor from the category $\mathsf{MD}$} build from the groups ${\mathsf{MD}}_{n}$ to the category $\mathsf{Mat}$ of matrices. In particular, such representations can be non-semisimple, pushing us into the world of indecomposable, rather than irreducible, constituent representations. Even for ${\mathbb{Z}}^{\binom{n}{2}}$ the classification problem for indecomposable representations is wild, so local representations are not easily decomposed.
(2.8) Here ‘local representation’ means in essence firstly that we consider representations $\varrho_{n}$ for a natural-number-indexed sequence of presented groups, and that in each presentation every generator is of the form $\alpha_{i}$ where $\alpha$ is taken from the same finite set, $i$ ranges over natural numbers, and every relation involving two non-adjacent $i$’s is a commutation. Secondly the representation $\varrho_{n}$ should also include in $\varrho_{n+1}$.
(2.9) It is clear from the relations $\mathfrak{R}$} that for $n>1$ the automorphism group of ${\mathsf{MD}}_{n}$ has a subgroup isomorphic to ${\mathbb{Z}}/2{\mathbb{Z}}\times{\mathbb{Z}}/2{\mathbb{Z}}$ generated by $s_{i}\leftrightarrow r_{i}$ and the simultaneous $(s_{i},r_{i})\leftrightarrow(s_{n-i},r_{n-i})$ for all $i$.
@@ 3.1. Warm-up: wreaths, their conjugacy classes, and their representations
(3.3) Recall further that there is a natural ‘bird-track’ calculus (cf. e.g. [7, 21]) for $G\wr\Sigma_{n}$ which decorates the strings of the usual diagram calculus for $\Sigma_{n}$ with elements of $G$. Writing $\tau\in G$ for a generating element (and $\sigma_{1}\in\Sigma_{n}$ as usual), then the calculus looks like this:
(all such figures here are taken directly from [21]). The calculus allows $\sigma_{1}\tau_{1}\sigma_{1}$ to be represented as a bead on the second track:
@@ 3.2. Generalised wreaths
(3.5) For $n\in\mathbb{N}$ consider the group ${\mathbb{Z}}^{{n}\choose{2}}$, with generators $x_{ij}$ for $1\leq i<j\leq n$. That is, ${\mathbb{Z}}^{{n}\choose{2}}=\;\langle x_{ij},i<j\;|x_{ij}x_{kl}=x_{kl}x_{ij}\rangle\;=\;\{\prod_{i<j}x_{ij}^{a_{ij}}\;|\;a_{ij}\in{\mathbb{Z}}\}$. We may write $g\in{\mathbb{Z}}^{{n}\choose{2}}$ as $(x_{12}^{a_{12}},x_{13}^{a_{13}},..)$. In particular that is $x_{12}=(x_{12}^{1},1,1,..,1)$ and $1=(1,1,..,1)$.
@@
(3.7) In ${\mathsf{MD}^{\prime}}_{n}$ define $\varsigma_{i}$ as $(1,\sigma_{i})=((1,1,\ldots,1),\sigma_{i})$ (where $\sigma_{i}$ is the elementary transposition in $\Sigma_{n}$) and $x_{ij}=(x_{ij},1)$.
Observe that ${\mathsf{MD}^{\prime}}_{n}$ admits a presentation with generators $x_{12}$ and $\varsigma_{i}$, $i=1,2,\ldots,n-1$ together with the usual symmetric group relations, the relation
@@ 3.3. The BaBeDa isomorphism
(3.8) Consider the map $\varphi$ on the generators of ${\mathsf{MD}^{\prime}}_{n}$ from (3.2) (and hence on the free group with these formal generators) to ${\mathsf{MD}}_{n}$ given by $\varsigma_{i}\mapsto s_{i}$ and $x_{12}\mapsto r_{1}s_{1}$. It is shown in [4, Prop.5.5] that $\varphi$ extends to an isomorphism ${\mathsf{MD}^{\prime}}_{n}\cong{\mathsf{MD}}_{n}$.
We call $\varphi$ the BaBeDa isomorphism (a portmanteau of the authors’ surnames). We also give the proof here to give a self-contained explanation of how the copy of ${\mathbb{Z}}^{{n}\choose{2}}$ arises. For this we need some preliminary results given next.
(3.9) The ‘virtual braid group’ ${\mathsf{VB}}_{n}$ is obtained by taking the same formal generators as $L_{n}$, and taking as relations the leftmost and rightmost relation in (1), as well as all relations in (2) and (3).
There is a map $\pi\colon{\mathsf{VB}}_{n}\to\Sigma_{n}$ given by letting both $r_{i}$ and $s_{i}$ map to the element $(i,i+1)$. This is clearly surjective, since the elementary transpositions $(i,i+1)$ generate $\Sigma_{n}$. Thus defining ${\mathsf{VP}}_{n}$ to be the kernel of this map $\pi$ gives that ${\mathsf{VB}}_{n}$ breaks down as a semi direct product ${\mathsf{VP}}_{n}\rtimes\Sigma_{n}$. It is proved in [3, Th.1] that ${\mathsf{VP}}_{n}$ admits a presentation with generators $\lambda_{kl}$, $1\leq k\neq l\leq n$, and relations
@@
$\lambda_{ki}\lambda_{kj}\lambda_{ij}=\lambda_{ij}\lambda_{kj}\lambda_{ki}$ (14)
where distinct letters represent distinct indices. The group ${\mathsf{VP}}_{n}$ is realised as a subgroup of ${\mathsf{VB}}_{n}$ by letting
@@
Observe for example that the first line forces $\pi(\lambda_{i,i+1})=1$ as required.
(3.10) It is also shown in [3, Le.1] that the subgroup ${\mathsf{VP}}_{n}\subset{\mathsf{VB}}_{n}$ satisfies the following conjugation relation
@@
First note that all relations involving only the symmetric group generators in each group (meaning the $s_{i}$s in ${\mathsf{MD}}_{n}$) are clearly preserved by both $\varphi$ and $\hat{\varphi}$.
@@
whenever $k\neq i$. This can be seen by observing that, using the conjugation rule (3.3) and that $\varphi(x_{ij})=\lambda_{ij}$,